How much Ione invest if she wants to have P19,640.62 at the end of 2.75 years at a bank that offers 11(1/4)% simple interest.

1
Expert's answer
2021-12-02T13:58:43-0500

"A=P(1+in).................(i)"

Where "A" is the accumulated amount,"P" is the principal invested ,"i" is the interest rate and "n" is the time period for accumulation

We are given:

"A= P19,640.62" , "i=11.25\%=0.1125" and "n=2.75" years

Substitute this values to equation "(i)" Above:

"A=P(1+in)"

"19,640.62 =P[1+(0.1125\times 2.75)]"

"\implies P={19640.62\over [1+(0.1125\times 2.75)]}=14999.9961813842\approx P15000"